Natalie, Chantal, and samara play together as a forward line in a hockey team. At the end of the season, Chantal had scored eight more goals than Natalie, while Samara had scored twice as many goals as Natalie. the three girls scored a total of 52 goals. How many did each girl score?
Answer by ikleyn(52776)

.
Natalie, Chantal, and samara play together as a forward line in a hockey team. At the end of the season,
Chantal had scored eight more goals than Natalie, while Samara had scored twice as many goals as Natalie.
the three girls scored a total of 52 goals. How many did each girl score?
~~~~~~~~~~


    Natalie + Chantal + Samara = 52   goals


      x     +  (x+8)  +  2x    = 52

               4x              = 52 - 8 = 44.


                x                       = 44/4 = 11.


Answer.  Natalie 11,  Chantal 19, Samara 22.

Solved.